Solve: 1 / x + 2 + 1 / x - 2 = 1 /x^2 - 4
a) -2
b) 2
c) 1/2
d) -1/2

Answer by mananth(12270)

Solve: 1 / x + 2 + 1 / x - 2 = 1 /x^2 - 4
LCM = (x+2)(x-2)
x-2 +x+2/ (x^2-4)= 1/(x^2-4)
x-2+x+2=1
2x=1
x=1/2